Solve by substitution
Y=x-3 and x +y=5

Answer by ikleyn(52787)

.
Solve by substitution
Y=x-3 and x +y=5
~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Hello,

you are lucky. 

They just made part of work for you by expressing y explicitly via x in the first equation.

So, simply substitute this expression y = x-3 into the second equation. You will get a single equation for x

x + (x-3) = 5,   or

2x - 3 = 5.

Hope you can easily complete this assignment.